# Predicting User Behavior in Smart Spaces with LLM-Enhanced Logs and Personalized Prompts

### v1

@@ Abstract
Enhancing the intelligence of smart systems, such as smart home, and smart vehicle, and smart grids, critically depends on developing sophisticated planning capabilities that can anticipate the next desired function based on historical interactions. While existing methods view user behaviors as sequential data and apply models like RNNs and Transformers to predict future actions, they often fail to incorporate domain knowledge and capture personalized user preferences. In this paper, we propose a novel approach that incorporates LLM-enhanced logs and personalized prompts. Our approach first constructs a graph that captures individual behavior preferences derived from their interaction histories. This graph effectively transforms into a soft continuous prompt that precedes the sequence of user behaviors. Then our approach leverages the vast general knowledge and robust reasoning capabilities of a pretrained LLM to enrich the oversimplified and incomplete log records. By enhancing these logs semantically, our approach better understands the user’s actions and intentions, especially for those rare events in the dataset. We evaluate the method across four real-world datasets from both smart vehicle and smart home settings. The findings validate the effectiveness of our LLM-enhanced description and personalized prompt, shedding light on potential ways to advance the intelligence of smart space.
Note: While this manuscript provides description of the data, we are not permitted to make these datasets publicly available due to restrictions imposed by the data provider.
journal: AAAI
[…]
organization=MOE KLINNS Lab, Xi’an Jiaotong University, city=Xi’an, postcode=710054, country=China
[…]
abstract
1 Introduction
@@
In today’s society, smart devices have become an indispensable part of our daily lives, significantly enriching our life experiences. However, despite the variety of functionalities these smart devices offer, they largely operate independently of one another, lacking deep inter-connectivity. This isolation hinders the formation of a truly “smart” systems. Take the smart vehicle scenario as an example: on a cold workday morning, a driver might open their car door, followed by the need to adjust the seat position, turn on the heating, set the navigation to plan a congestion-free route, and play their favorite music. While these functions are straightforward, manually executing each one can be tedious. Enabling a simple action, like opening a door, to automatically trigger all related tasks would significantly enhance the user experience in smart settings. Simplifying the interaction between users and smart devices to allow the devices to understand and automatically adapt to users’ needs in various scenarios represents an urgent yet challenging goal [1].
@@
This vision intersects with two main areas of research: human activity prediction and sequential recommendation. In human activity prediction, previous works in the context of smart space attempt to build models from historical interaction data between users and their environments, with the goal of automating the recognition of users’ intentions and plan the next move [2, 3, 4, 5]. For instance, deep learning models have been employed to analyze various sensor data within smart homes to forecast subsequent user activities [6]. On the other hand, sequential recommendation focuses on the online activity context suggesting content that users may find engaging based on their behaviors [7, 8]. This includes analyzing sequences of user online interactions—such as movie watching, web browsing, clicking, bookmarking, and purchasing—to tailor recommendations of potentially interesting content like movies or products.
@@
However, in real-world smart spaces, existing methods face significant challenges that differ markedly from those in recommendation systems. While recommendation systems can access extensive side information—including the names, brands, and attributes of purchased items [8], and content of movies that users are interested in—data logs from devices and sensors in smart spaces are often sparse or even incomplete. Consequently, earlier methods heavily often employ one-hot encoding to represent different events as unrelated IDs, missing the inherent connections between related actions. For example, opening a refrigerator and preheating an oven are both linked to the intention of eating, yet are treated as unrelated by one-hot encoding. This approach fails to recognize the natural associations between activities, leading to a disconnect in representations, particularly for low-frequency actions that are poorly represented.
@@
Additionally, due to the limited number of users in previous smart space datasets [2, 9, 10, 5], prior research in smart spaces typically employs a single model to predict a range of user actions, neglecting the individual preferences that result in behavioral differences. For instance, the routine of going to sleep could involve actions like drawing the curtains and switching off the lights for some, while for others, it might include playing soft music and setting an alarm. This variation highlights the inadequacy of a one-size-fits-all approach in predicting user actions without considering individual preferences, which is crucial for effective smart space automation.
@@
To address the challenges in predicting user behavior within smart systems, we propose a novel approach that incorporates LLM-enhanced logs and personalized prompts. This method taps into the vast general knowledge and robust reasoning capabilities of a pretrained LLM, using in-context learning to enrich the oversimplified and incomplete log records. By enhancing these logs semantically, we can better understand the user’s actions and intentions. Additionally, our approach constructs a graph that captures individual behavior preferences derived from their interaction histories. This graph effectively transforms into a soft continuous prompt that precedes the sequence of user behaviors. Through prompt tuning, the model is guided to make differentiated predictions about user behavior based on this prompt.
@@
We primarily evaluate our approach through experiments conducted on two large-scale, real-world dataset of user interactions with smart vehicles, complemented by two additional datasets focusing on smart home settings. The experimental outcomes demonstrate the superior performance of our approach when compared to state-of-the-art methods in terms of predicting a user’s next move within smart vehicle contexts. By incorporating LLMs for description augmentation, our approach significantly boosts performance in scenarios involving low-frequency events. Furthermore, our framework demonstrates robust adaptability, effectively handling cold-start scenarios with new users and seamlessly transitioning to smart home contexts. The contributions are summarized as follows:
  • 1.

### v2

@@ Abstract
Enhancing the intelligence of smart systems, such as smart homes, smart vehicles, and smart grids, critically depends on developing sophisticated planning capabilities that can anticipate the next desired function based on historical interactions. While existing methods view user behaviors as sequential data and apply models like RNNs and Transformers to predict future actions, they often fail to incorporate domain knowledge and capture personalized user preferences. In this paper, we propose a novel approach that incorporates LLM-enhanced logs and personalized prompts. Our approach first constructs a graph that captures individual behavior preferences derived from their interaction histories. This graph effectively transforms into a soft continuous prompt that precedes the sequence of user behaviors. Then our approach leverages the vast general knowledge and robust reasoning capabilities of a pretrained LLM to enrich the oversimplified and incomplete log records. By enhancing these logs semantically, our approach better understands the user’s actions and intentions, especially for those rare events in the dataset. We evaluate the method across four real-world datasets from both smart vehicle and smart home settings. The findings validate the effectiveness of our LLM-enhanced description and personalized prompt, shedding light on potential ways to advance the intelligence of smart space.
Introduction
@@
In today’s society, smart devices have become an indispensable part of our daily lives, significantly enriching our life experiences. However, despite the variety of functionalities these smart devices offer, they largely operate independently of one another, lacking deep inter-connectivity. This isolation hinders the formation of truly “smart” systems. Take the smart vehicle scenario as an example: on a cold workday morning, a driver might open their car door, followed by the need to adjust the seat position, turn on the heating, set the navigation to plan a congestion-free route, and play their favorite music. While these functions are straightforward, manually executing each one can be tedious. Enabling a simple action, like opening a door, to automatically trigger all related tasks would significantly enhance the user experience in smart settings. Simplifying the interaction between users and smart devices to allow the devices to understand and automatically adapt to users’ needs in various scenarios represents an urgent yet challenging goal (Desolda, Ardito, and Matera 2017).
@@
This vision intersects with two main areas of research: human activity prediction and sequential recommendation. In human activity prediction, previous works in the context of smart space attempt to build models from historical interaction data between users and their environments, with the goal of automating the recognition of users’ intentions and planning the next move (Tax 2018; Maharjan et al. 2019; Mohamed, El-Kilany, and El-Tazi 2022; Dunne, Morris, and Harper 2023). For instance, deep learning models have been employed to analyze various sensor data within smart homes to forecast subsequent user activities (Krishna et al. 2018). On the other hand, sequential recommendation focuses on the online activity context suggesting content that users may find engaging based on their behaviors (Hou et al. 2022; Li et al. 2023). This includes analyzing sequences of user online interactions, such as movie watching, web browsing, clicking, bookmarking, and purchasing, to tailor recommendations of potentially interesting content, like movies or products.
@@
However, in real-world smart spaces, existing methods face challenges that differ markedly from those in recommendation systems. While recommendation systems can access extensive side information, which includes the names, brands, and attributes of purchased items (Li et al. 2023), and the content of movies that interest users, data logs from devices and sensors in smart spaces are often sparse or even incomplete. Consequently, earlier methods that employed one-hot encoding to represent different events as unrelated IDs frequently missed the inherent connections between related actions. For example, opening a refrigerator and preheating an oven are both linked to the intention of eating, yet are treated as unrelated by one-hot encoding. This approach fails to recognize the natural associations between activities, leading to a disconnect in representations, particularly for low-frequency actions that are poorly represented.
@@
Additionally, due to the limited number of users in previous smart space datasets (Tax 2018; Park et al. 2022; Wang et al. 2022b; Dunne, Morris, and Harper 2023), prior research in smart spaces typically employs a single model to predict a range of user actions, neglecting the individual preferences that result in behavioral differences. For instance, the routine of going to sleep could involve actions like drawing the curtains and switching off the lights for some, while for others, it might include playing soft music and setting an alarm. This variation highlights the inadequacy of a one-size-fits-all approach in predicting user actions without considering individual preferences, which is crucial for effective smart space automation.
@@
To address the challenges in predicting user behavior within smart spaces, we propose a novel approach that incorporates LLM-enhanced logs and personalized prompts. This method taps into the vast general knowledge and robust reasoning capabilities of a pretrained LLM, using in-context learning to enrich the oversimplified and incomplete log records. By enhancing these logs semantically, we can better understand the user’s actions and intentions. Additionally, our approach constructs a graph that captures individual behavior preferences derived from their interaction histories. This graph effectively transforms into a soft continuous prompt that precedes the sequence of user behaviors. Through prompt tuning, the model is guided to make varied predictions about user behavior based on this prompt.
@@
We primarily evaluate our approach through experiments conducted on two large-scale, real-world datasets of user interactions with smart vehicles, complemented by two additional datasets focusing on smart home settings. The experimental outcomes demonstrate the superior performance of our approach when compared to state-of-the-art methods in terms of predicting a user’s next move within smart vehicle contexts. By incorporating LLMs for description augmentation, our approach significantly boosts performance in scenarios involving low-frequency events. Furthermore, our framework demonstrates robust adaptability, effectively handling cold-start scenarios with new users and seamlessly transitioning to smart home contexts. The contributions are summarized as follows: